# Preprint Notice

### v3

@@ III-B3 Practical Implications and Real-World Consequences
III-C Dedicated Private Lines Are Essential for High-Speed Testing
@@
From 2015-2019, the zx R&D was carried out on a testbed at SLAC National Accelerator Laboratory, using a 5,000-mile, 100 Gbps loop provisioned by ESnet. After 2019, collaboration with ESnet continued [34]. Nevertheless, accessing their 100G Software Defined Network (SDN) Testbed required a formal application process [10], hindering agile development. This need for a fully controlled, on-demand test environment led us to evaluate commercial WAN simulators, which we found to be prohibitively expensive and functionally limited.
A turning point was the discovery of a 2012 technical report, “Validating Linux Network Emulation” [35], which outlined a method using Linux's built-in traffic control (tc [57]) and network emulation (tc-netem [56]) tools. Recognizing the potential of this software-based approach to create a high-fidelity, 100 Gbps capable testbed at a fraction of the cost, we implemented an enhanced version. This version extended functionality and automation beyond the original method. The resulting testbed (Fig. 7) —built at the former Intel Swindon Lab in collaboration with Intel and HPE—consists of two HPE DL380 Gen 11 servers (configured as the appliances under test) connected via a switch to two Intel "latency servers." These servers use tc and tc-netem to impose precise delay, accurately simulating transcontinental links in an automated, remotely manageable setup.
Figure 7: Intel Corp. arranged its former Swindon Lab in Swindon, U.K., to collaborate with Zettar using an enhanced version of the approach created by Dr. Ezra Kissel. The essential components are labeled explicitly.
Figure 8: Bulk transfer sweeps vs three simulated latency values: 10 ms, 50 ms, and 100 ms, corresponding respectively to Los Angeles to San Francisco, California, U.S.; Singapore to Alaska, U.S.; and Singapore to Atlanta, Georgia, U.S.
@@
Figure 9: Streaming transfer sweeps vs three simulated latency values, 10ms, 50ms, and 100ms. Note that the data rate levels are quite close to their counterparts from the bulk transfer sweeps shown in Fig. 6.
  • Systems Focus: Unlike commercial WAN simulators or network testbeds that isolate network performance (often using Random Access Memory (RAM)-to-RAM methods), this platform is uniquely suited to measure true burst-buffer to burst-buffer data transfer rates. This is essential because, as established in Section III-A, persistent I/O is the dominant constraint for petascale transfers.
@@
The full testbed benefits and implementation details, together with expanding it to speed $\geq 100$ Gbps, warrant a separate, in-depth tutorial. See also Section VII.
III-D Increasing network bandwidth increases transfer rates
@@
Figure 10: For production storage to support moving data at scale and speed well, it must have high enough throughput and low latency.
@@
III-E Powerful CPUs are essential for high transfer rates
A common consensus in high-performance computing is that powerful CPUs are essential for high transfer rates. During 2023, while discussing with our industry collaborators regarding CPU selection for data movement appliances, the prevailing feedback was to select high-end, high-core-count models. We held reservations about this approach. Furthermore, we also had skepticism about suggestions to utilize CPUs with embedded hardware acceleration such as Quick Assist Technology (QAT) [30], based on the observation that such acceleration can tie software to specific vendor hardware and drivers, increasing complexity and potential "software bloat."
This skepticism was grounded in years of deployment experience, which demonstrated that zx consistently achieves high performance without requiring a high core count; typically, 12-24 cores are sufficient. While resources like ESnet's fasterdata recommend high clock rates for good encryption performance [11], our observations suggested an alternative path. We hypothesized that fewer cores with CPU-built-in encryption acceleration and moderate clock rates would reduce context switches (improving software efficiency) and lower energy consumption, thereby reducing the TCO. This led to the selection of the Intel Xeon 5418N, a mid-range model [29].
@@ III-E Powerful CPUs are essential for high transfer rates
III-F Virtualization and cloud are universally useful
We recently received an inquiry about evaluating zx for a site in South Africa needing to transfer data to Europe at higher rates. The site's proposal to use a virtual machine (VM) immediately raised concerns, illustrating a fundamental mismatch between virtualized environments and data movement performance.
Virtualization technologies have been a mainstay for over two decades [46]. A critical determinant of their performance for data-intensive workloads is the level of system control accessible to the user, summarized in Table VI.
TABLE VI: Linux Kernel Parameter Scope: Virtual Machine vs. Container.
Setting location Target Platform Scope of sysctrl Change
Inside VM (Guest OS) Virtual Machine (e.g., KVM) Affects only the Guest OS kernel
Inside Container (default) Container (Docker/Podman) Affects only a small, namespaced subset of settings for the container. Cannot change host-level settings.
Inside Container (privileged) Container (Docker/Podman) Can change host-level settings, affecting both the host and all running containers.
On Host VM Host or Container Host Affects the underlying physical server and, in the case of containers, the default for all running containers.
@@ III-F Virtualization and cloud are universally useful
    Interrupt Overhead: Even with Single Root I/O Virtualization (SR-IOV) [41], interrupt handling in VMs incurs measurable overhead due to interception and remapping by the hypervisor.
@@
    Host-Level Dependency: If the host kernel is not tuned to mitigate latency (as established in Section III-A), adjustments within the guest OS are largely ineffective.
Achieving consistent, latency-insensitive data transfer performance within a VM is intrinsically constrained by the architecture. Cloud providers inherit these limitations. Coordinating the kernel parameter adjustments necessary for maximum VM performance is application-dependent (see Table III) and practically infeasible for a heterogeneous customer base.
III-F1 The Architectural Cost of Cloud Abstraction
@@
Standard cloud workarounds, such as Amazon Web Services (AWS) multipart uploads [3], fragment a simple file transfer into a multi-step coordination process—effectively a highly fragmented workflow [2] for data movement.
III-F2 A Critical Examination of Cloud Provider Metrics
The prevailing paradigm of general-purpose cloud data paths, regardless of provider, introduces inherent architectural conflicts for high-performance data movement workflows. This challenge becomes acute at the petascale level. Analysis of major U.S. cloud providers reveals a focus on theoretical network bandwidth, often decoupled from storage I/O and end-to-end performance. The info in Tables VII-IX, extracted from official provider documentation, illustrates this focus.
@@ III-F2 A Critical Examination of Cloud Provider Metrics
Maximum network bandwidth figures—such as 1,000 Gbps or 400 Gbps —are often theoretically impressive but unrealizable in practice, as storage subsystems cannot always sustain the throughput. The published info from three leading hyperscalers, shown in Tables X should make this aspect more clear.
TABLE VII: Google Cloud Platform (GCP) virtual instance parameters.
@@
TABLE VIII: AWS virtual instance parameters.
@@
TABLE IX: Azure virtual instance parameters.
@@
  • Data is produced predominantly at the edge (e.g., LCLS-II [53]).
@@
  • Treating this latency [59] and [37] deserves scrutiny.
TABLE X: Typical performance information published by hyperscalers.
Cloud Provider Workload Quantitative Metric Throughput Implication
AWS Western Digital HPC Simulation 2.5 million tasks on 1 million vCPU cluster in 8 hours Massive-scale, low-latency, high-bandwidth communication; reduces job time from 20 days on-prem to 8 hours in cloud.
GCP PGS Seismic Processing 202k on-prem cores → 1.2M vCPUs in cloud 6× increase in compute/data-transfer; among world’s largest supercomputers if continuous.
Azure University of Bath HPC 250 on-prem nodes → thousands in Azure Provides on-demand, high-speed burst capacity; ultra-high data access elasticity.
@@
III-F3 The Antidote: A Co-Designed Data Path for the Cloud
The solution is not to work within the existing cloud models’ constraints, but to architect a parallel data path coexisting with it. By embedding zx in DPUs [42, 62], we create a high-performance path that bypasses the cloud software stack. This provides:
@@ III-F3 The Antidote: A Co-Designed Data Path for the Cloud
III-F4 Empirical Evidence of the Cloud Data Movement Limitations
The performance penalties imposed by standard cloud data paths are measurable. Using the GoToCloud platform on AWS [38], a comparative analysis conducted by High Energy Accelerator Research Organization (KEK) [27] vividly illustrates this point. As summarized in Fig. 11, the transfer of a 1.2 TiB Cryo-electron microscopy (Cryo-EM) dataset was measured with zx vs. aws-cli
Figure 11: Transfer comparison of a 1.2 TiB Cryo-EM dataset from KEK to AWS Regions using aws-cli and zx.
@@ Results:
IV Methodology and Reproducibility
To enable independent validation of the performance claims presented in this paper, we have provided all essential configuration details required to replicate our testing environment. This includes the exact kernel parameters, latency simulation values, and visualization commands, all of which are publicly available in a GitHub repository [19]. The iperf3 benchmark tool used in comparative analysis is also publicly available from its official repositories [36].
@@ IV Methodology and Reproducibility
V Deployments
V-A Validation Platforms
This testing is part of a decade-long, sustained effort to validate the co-design principle across diverse production environments, as documented in prior studies [53, 14, 15, 28, 40, 17, 21].
@@ V-A Validation Platforms
  • Two mini appliances deployed at two distinct sites in Taiwan, connected via the Taiwan Advanced Research and Education Network (TWAREN) [52].
V-B Data Availability and Licensing
In the spirit of reproducible research, all elements required to replicate the testing and benchmarking methodology presented in this paper are provided in a public GitHub repository [19]. The zx software data mover is a commercially licensed product and is not available for public distribution.
VI Conclusions
The arguments presented in this paper against over-reliance on isolated, formulaic optimizations are rooted in a system perspective. As in the design of composite laminates [20], where macroscopic properties of a composite plate result from the arrangement of discrete piles, the performance of a data movement system emerges from the interplay of discrete, non-linear components: CPU cores, memory hierarchies, storage devices, network interfaces, among others.
The holistic co-design principle central to this work represents the evolution of a decade-plus sustained and continuing engineering effort. This lineage traces back to foundational architectural blueprints developed at SLAC National Accelerator Laboratory [17] and subsequently validated through formal DOE/ESnet evaluation [34]. The architecture’s efficacy has been demonstrated through high-stakes production trials at the Linac Coherent Light Source (LCLS) [53] and through extensive public discourse [21], [18], [22]. This maturing perspective has recently been extended to modern hardware platforms, demonstrating that consistent, high-performance data movement can be seamlessly carried out on Data Processing Units (DPUs) [23], [62]. This maturing perspective underscores that consistent, high-performance data movement has always been fundamentally a systems engineering challenge rather than an isolated network optimization problem.
@@ VI Conclusions
The scalability of this architectural approach was demonstrated as early as 2018 in the petabyte-transfer setup [14] and [15] shown in Fig. 12 (reproduced from [34] Appendix 6.1). This scale-out design used multiple older servers with zx-aggregated 10Gbps interfaces, proving that the invariant software architecture could distribute workload across nodes without cluster management overhead. This consistency ensures that the logical appliance design—whether implemented as a 10Gbps mini-appliance or a 400Gbps core node—remains uniform across scales. For example, modern 2U2N servers, e.g., the SYS- [^2] BT-DNR [51] (Fig. 13), enable the same scale-out efficiency within a single chassis. It extends our testbed-based methodology to 400 Gbps and beyond, using the same architectural principles that successfully moved petabytes in 2018.
Figure 12: Zettar’s testbed at SLAC from 2015 to 2019. See page 19 of [40] for more details of the configuration. (Copyright Zettar Inc. Reproduced with permission for this publication.)
Figure 13: The Supermicro 2U 2-Node BigTwin with 12 hot-swap 2.5" NVMe drives per node (Diagram by Chin Fang).

### v4

@@ 3.2.3. Practical Implications and Real-World Consequences
3.3. Dedicated Private Lines Are Essential for High-Speed Testing
@@
From 2015-2019, the zx R&D was carried out on a testbed at SLAC National Accelerator Laboratory, using a 5,000-mile, 100 Gbps loop provisioned by ESnet. After 2019, collaboration with ESnet continued (Kissel and Fang, 2020c). Nevertheless, accessing their 100G Software Defined Network (SDN) Testbed required a formal application process (Energy Sciences Network (2025f), ESnet), hindering agile development. This need for a fully controlled, on-demand test environment led us to evaluate commercial WAN simulators, which we found to be prohibitively expensive and functionally limited.
A turning point was the discovery of a 2012 technical report, “Validating Linux Network Emulation” (Kissel and Swany, 2012), which outlined a method using Linux's built-in traffic control (tc (The Linux man-pages project, 2025b)) and network emulation (tc-netem (The Linux man-pages project, 2025a)) tools. Recognizing the potential of this software-based approach to create a high-fidelity, 100 Gbps capable testbed at a fraction of the cost, we implemented an enhanced version. This version extended functionality and automation beyond the original method. The resulting testbed (Fig. 7) —built at the former Intel Swindon Lab in collaboration with Intel and HPE—consists of two HPE DL380 Gen 11 servers (configured as the appliances under test) connected via a switch to two Intel "latency servers." These servers use tc and tc-netem to impose precise delay, accurately simulating transcontinental links in an automated, remotely manageable setup.
Figure 7. Intel Corp. arranged its former Swindon Lab in Swindon, U.K., to collaborate with Zettar using an enhanced version of the approach created by Dr. Ezra Kissel. The essential components are labeled explicitly.
The topology of the testbed setup at Intel Swindon Lab showing two HPE servers connected via latency simulation servers.
Figure 8. Bulk transfer sweeps vs three simulated latency values: 10 ms, 50 ms, and 100 ms, corresponding respectively to Los Angeles to San Francisco, California, U.S.; Singapore to Alaska, U.S.; and Singapore to Atlanta, Georgia, U.S.
Graph showing streaming transfer performance vs. simulated latency values (10ms, 50ms, 100ms) on the 100 Gbps-capable testbed.
@@
Figure 9. Streaming transfer sweeps vs three simulated latency values, 10ms, 50ms, and 100ms. Note that the data rate levels are quite close to their counterparts from the bulk transfer sweeps shown in Fig. 6.
Graph showing streaming transfer performance vs. simulated latency values (10ms, 50ms, 100ms) on the aforementioned 100 Gbps-capable testbed.
  • Systems Focus: Unlike commercial WAN simulators or network testbeds that isolate network performance (often using Random Access Memory (RAM)-to-RAM methods), this platform is uniquely suited to measure true burst-buffer to burst-buffer data transfer rates. This is essential because, as established in Section 3.1, persistent I/O is the dominant constraint for petascale transfers.
@@
The full testbed benefits and implementation details, together with expanding it to speed $\geq 100$ Gbps, warrant a separate, in-depth tutorial. See also Section 7.
3.4. Increasing network bandwidth increases transfer rates
@@
Figure 10. For production storage to support moving data at scale and speed well, it must have high enough throughput and low latency.
A diagram illustrating storage throughput and latency requirements needed to support data transfer at speed and scale.
@@
3.5. Powerful CPUs are essential for high transfer rates
A common consensus in high-performance computing is that powerful CPUs are essential for high transfer rates. During 2023, while discussing with our industry collaborators regarding CPU selection for data movement appliances, the prevailing feedback was to select high-end, high-core-count models. We held reservations about this approach. Furthermore, we also had skepticism about suggestions to utilize CPUs with embedded hardware acceleration such as Quick Assist Technology (QAT) (Intel, 2025b), based on the observation that such acceleration can tie software to specific vendor hardware and drivers, increasing complexity and potential "software bloat."
This skepticism was grounded in years of deployment experience, which revealed that zx consistently achieves high performance without requiring a high core count; typically, 12-24 cores are sufficient. While resources like ESnet's fasterdata recommend high clock rates for good encryption performance (Energy Sciences Network (2025c), ESnet), our observations suggested an alternative path. We hypothesized that fewer cores with CPU-built-in encryption acceleration and moderate clock rates would reduce context switches (improving software efficiency) and lower energy consumption, thereby reducing the TCO. This led to the selection of the Intel Xeon 5418N, a mid-range model (Intel, 2025a).
@@ 3.5. Powerful CPUs are essential for high transfer rates
3.6. Virtualization and cloud are universally useful
We recently received an inquiry about evaluating zx for a site in South Africa needing to transfer data to Europe at higher rates. The site's proposal to use a virtual machine (VM) immediately raised concerns, as it revealed a common misconception: that virtualized environments can deliver the performance required for high-speed data movement.
Virtualization technologies have been a mainstay for over two decades (Red Hat, 2025c). Nevertheless, a critical determinant of their performance for intensive data movement workloads is the level of system control accessible to the user. Table 6 summarizes the scope of Linux kernel parameters across virtual machines and containers.
Table 6. Linux Kernel Parameter Scope: Virtual Machine vs. Container
Location Platform Scope
Inside VM VM (e.g. KVM) Affects only the guest OS kernel.
Container Docker/Podman Affects namespaced subset; no host change.
Privileged Docker/Podman Can change host settings, affecting all.
On Host Physical host Affects physical server and all VMs and containers.
@@ 3.6. Virtualization and cloud are universally useful
    Interrupt Overhead: Even with Single Root I/O Virtualization (SR-IOV) (NVIDIA, 2025a), interrupt handling in VMs incurs measurable overhead due to interception and remapping by the hypervisor.
@@
    Host-Level Dependency: If the host kernel is not tuned to mitigate latency (as established in Section 3.1), adjustments within the guest OS are largely ineffective.
Achieving consistent, latency-insensitive data transfer performance within a VM is intrinsically constrained by the architecture. Cloud providers inherit these limitations. Coordinating the kernel parameter adjustments necessary for maximum VM performance is application-dependent (see Table 3) and practically infeasible for a heterogeneous customer base.
3.6.1. The Architectural Cost of Cloud Abstraction
@@
Standard cloud workarounds, such as Amazon Web Services (AWS) multipart uploads (Amazon Web Services, 2025), fragment a simple file transfer into a multi-step coordination process—effectively a highly fragmented workflow (Acharya and Sirinterlikci, 2010) for data movement.
3.6.2. A Critical Examination of Cloud Provider Metrics
The prevailing paradigm of general-purpose cloud data paths, regardless of provider, introduces inherent architectural conflicts for high-performance data movement workflows. This challenge becomes acute at the petascale level. Analysis of major U.S. cloud providers reveals a focus on theoretical network bandwidth, often decoupled from storage I/O and end-to-end performance. The info in Tables 7-9, extracted from official provider documentation, illustrates this focus.
@@ 3.6.2. A Critical Examination of Cloud Provider Metrics
Maximum network bandwidth figures—such as 1,000 Gbps or 400 Gbps —are often theoretically impressive but unrealizable in practice, as storage subsystems cannot always sustain the throughput. The published info from three leading hyperscalers, shown in Tables 10 should make this aspect more clear.
Table 7. Google Cloud Platform (GCP) virtual instance parameters.
@@
Table 8. AWS virtual instance parameters.
@@
Table 9. Azure virtual instance parameters.
@@
  • Data is produced predominantly at the edge (e.g., LCLS-II (Thayer and Fang, 2019)).
@@
  • Treating this latency (WSJ Staff, 2025) and (Morales, 2025) deserves scrutiny.
Table 10. Typical performance information published by hyperscalers.
Cloud Provider Workload Metric Throughput Implication
AWS Western Digital HPC Simulation 2.5M tasks on 1M vCPU cluster in 8 hours Massive-scale, low-latency, high-bandwidth communication; reduces job time from 20 days on-prem to 8 hours in cloud.
[…]
Azure Univ. of Bath HPC 250 on-prem nodes $\rightarrow$ thousands in Azure Provides on-demand, high-speed burst capacity; ultra-high data access elasticity.
@@
3.6.3. The Antidote: A Co-Designed Data Path for the Cloud
The solution is not to work within the existing cloud models’ constraints, but to architect a parallel data path coexisting with it. By embedding zx in DPUs (NVIDIA, 2025b; Zettar Inc., 2024), we create a high-performance path that bypasses the cloud software stack. This provides:
@@ 3.6.3. The Antidote: A Co-Designed Data Path for the Cloud
3.6.4. Empirical Evidence of the Cloud Data Movement Limitations
The performance penalties imposed by standard cloud data paths are measurable. Using the GoToCloud platform on AWS (Moriya et al., 2024), a comparative analysis conducted by High Energy Accelerator Research Organization (KEK) (High Energy Accelerator Research Organization (2025), KEK) vividly illustrates this point. As summarized in Fig. 11, the transfer of a 1.2 TiB Cryo-electron microscopy (Cryo-EM) dataset was measured with zx vs. aws-cli
Figure 11. Transfer comparison of a 1.2 TiB Cryo-EM dataset from KEK to AWS Regions using aws-cli and zx.
A bar chart comparing transfer performance of zx vs. aws-cli for Cryo-EM dataset transfers from KEK to AWS Tokyo and Virginia.
@@ Results:
4. Methodology and Reproducibility
To enable independent validation of the performance claims presented in this paper, we have provided all essential configuration details required to replicate our testing environment. This includes the exact kernel parameters, latency simulation values, and visualization commands, all of which are publicly available in a GitHub repository (Fang, 2025). The iperf3 benchmark tool used in comparative analysis is also publicly available from its official repositories (man7.org, 2025).
@@ 4. Methodology and Reproducibility
5. Deployments
5.1. Validation Platforms
This testing is part of a decade-long, sustained effort to validate the co-design principle across diverse production environments, as documented in prior studies (Thayer and Fang, 2019; Energy Sciences Network (2017), ESnet; Energy Sciences Network (2018), ESnet; ICM, 2019; NSCC, 2019; Fang et al., 2016; Fang, 2019).
@@ 5.1. Validation Platforms
  • Two mini appliances deployed at two distinct sites in Taiwan, connected via the Taiwan Advanced Research and Education Network (TWAREN) (TaiWan Advanced Research and Education Network (2025), TWAREN).
  • Independent validation conducted at KEK in Japan using the GoToCloud platform on AWS, as reported in Section 3.6.4.
5.2. Data Availability and Licensing
In the spirit of reproducible research, all elements required to replicate the testing and benchmarking methodology presented in this paper are provided in a public GitHub repository (Fang, 2025). The zx software data mover is a commercially licensed product and is not available for public distribution.
6. Conclusions
The arguments presented in this paper against over-reliance on isolated, formulaic optimizations are rooted in a system perspective. As in the design of composite laminates (Fang and Springer, 1993), where macroscopic properties of a composite plate result from the arrangement of discrete piles, the performance of a data movement system emerges from the interplay of discrete, non-linear components: CPU cores, memory hierarchies, storage devices, network interfaces, among others.
The holistic co-design principle central to this work represents the evolution of a decade-plus sustained and continuing engineering effort. This lineage traces back to foundational architectural blueprints developed at SLAC National Accelerator Laboratory (Fang et al., 2016) and subsequently validated through formal DOE/ESnet evaluation (Kissel and Fang, 2020c). The architecture’s efficacy has been demonstrated through high-stakes production trials at the Linac Coherent Light Source (LCLS) (Thayer and Fang, 2019) and through extensive public discourse (Fang, 2019), (Fang and Kissel, 2021), (Fang, 2021). This maturing perspective has recently been extended to modern hardware platforms, demonstrating that consistent, high-performance data movement can be seamlessly carried out on Data Processing Units (DPUs) (Fang, 2022), (Zettar Inc., 2024). This maturing perspective underscores that consistent, high-performance data movement has always been fundamentally a systems engineering challenge rather than an isolated network optimization problem.
@@ 6. Conclusions
The scalability of this architectural approach was demonstrated as early as 2018 in the petabyte-transfer setup (Energy Sciences Network (2017), ESnet) and (Energy Sciences Network (2018), ESnet) shown in Fig. 12 (reproduced from (Kissel and Fang, 2020c) Appendix 6.1). This scale-out design used multiple older servers with zx-aggregated 10 Gbps interfaces, proving that the invariant software architecture could distribute workload across nodes without cluster management overhead. This consistency ensures that the logical appliance design—whether implemented as a 10 Gbps mini-appliance or a 400 Gbps core node—remains uniform across scales. For example, modern 2U2N servers, e.g., the SYS- [^2] BT-DNR (Supermicro, 2025) (Fig. 13), enable the same scale-out efficiency within a single chassis. It extends our testbed-based methodology to 400 Gbps and beyond, using the same architectural principles that successfully moved petabytes in 2018.
Figure 12. Zettar’s testbed at SLAC from 2015 to 2019. See page 19 of (NSCC, 2019) for more details of the configuration. (Copyright Zettar Inc. Reproduced with permission for this publication.)
Photograph of Zettar’s testbed at SLAC National Accelerator Laboratory from 2015-2019 showing server rack setup.
Figure 13. The Supermicro 2U 2-Node BigTwin with 12 hot-swap 2.5" NVMe drives per node (Diagram by Chin Fang).
Photograph of Supermicro 2U 2-Node BigTwin server with 12 hot-swap 2.5" NVMe drives per node.